OpenStudy (anonymous):

Write the equation of the line with an undefined slope, passing through the point (2, 5)

OpenStudy (alexwee123):

x=2

OpenStudy (anonymous):

thnk u

OpenStudy (kropot72):

The standard form of the equation is y = mx + b Substituting the given values for x and y gives: 5 = 2m + b b = 5 - 2m therefore the equation of the line with undefined slope m is as follows: y = mx + (5 - 2m)